# A quantitative analysis of semantic information in deep representations of text and images

### v3

@@ Abstract
Deep neural networks are known to develop similar representations for semantically related data, even when they belong to different domains, such as an image and its description, or the same text in different languages. We present a method for quantitatively investigating this phenomenon by measuring the relative information content of the representations of semantically related data and probing how it is encoded into multiple tokens of large language models (LLMs) and vision transformers. Looking first at how LLMs process pairs of translated sentences, we identify inner “semantic” layers containing the most language-transferable information. We find moreover that, on these layers, a larger LLM (DeepSeek-V3) extracts significantly more general information than a smaller one (Llama3.1-8B). Semantic information of English text is spread across many tokens and it is characterized by long-distance correlations between tokens and by a causal left-to-right (i.e., past-future) asymmetry. We also identify layers encoding semantic information within visual transformers. We show that caption representations in the semantic layers of LLMs predict visual representations of the corresponding images. We observe significant and model-dependent information asymmetries between image and text representations.
@@ 1 Introduction
Current research explores the conjecture that Large Language Models (LLMs) possess a universal language of thought, enabling them to process concepts across different languages peng2024conceptspacealignmentmultilingual ; anthropic_claude ; platonic_hypothesis ; singh-2019-bert ; brinkmann2025largelanguagemodelsshare . A compelling example comes from Anthropic’s study of the Claude model anthropic_claude , which shows that the network occasionally reasons within a shared conceptual space across languages. Intriguingly, the extent of this cross-lingual alignment appears to grow with model size, suggesting that universality may be an emergent property of scale. Ref. platonic_hypothesis further argues more generally that representations in deep networks, including but not limited to LLMs, are converging, potentially pointing toward shared underlying conceptual structures.
@@
In this work, we seek to turn the qualitative notion that deep networks develop semantically meaningful shared representations into quantitative measures, identifying where in the network, and to what extent, semantic information is encoded. We can cast the identification of shared semantic content as a problem of relative information: different networks learn distributions defined on different data domains, such as images and captions. Those distributions loosely span a common underlying subspace, associated with shared semantic information. The support of the different distributions does not reduce to this subspace, but spans other subspaces, which are data and architecture specific. To quantify relative information we need a similarity criterion that is both (i) asymmetric, as there may be a partial order relation between models and representations (qualitatively, the subspaces which are data specific might have different dimensions); (ii) computationally efficient for representations of dimension of order $10^{5}-10^{6}$, which is the typical number of neurons in a deep representation. Cross-entropy measures how difficult it is to encode an event from a distribution into another and would be the ideal choice, but estimating it is computationally difficult due to the dimension of the representations. On the other hand, local methods like the average number of nearest neighborsdoimo2020 ; platonic_hypothesis fail to capture the asymmetry between spaces. To overcome the limitations, we employ the information imbalance II : a method that leverages conditional ranking to provide an asymmetric measure of relative mutual predictivity, which has been shown to be an excellent proxy of the cross entropy Causality-II .
@@
We analyze the representation of translations of the same sentences in different languages, using one of the most powerful open-source language models, DeepSeek-V3 with 671B parameters deepseekv3 , comparing it to the mid-sized Llama3.1-8b LLM Meta:2024 . Moreover, we analyze the deep representations of pairs of images and of human-generated descriptions, and of pairs of images depicting the same object, with the goal of capturing the alignment between image and caption representations in different architectures and the relative predictive power of images and texts. As visual models, we employ image-gpt-large pmlr-v119-chen20s and DinoV2-large oquabDINOv2LearningRobust2024 .
@@
  • We analyze the DeepSeek-V3 language model and measure, layer by layer, the relative information content between representations of sentence pairs that are translations of each other. This reveals a broad region of the network, robust across language pairs, that consistently encodes shared semantic content, and whose extent increases with model size.
  • We show that in deep layers, information is distributed across tokens through long-range correlations. Moreover, we observe that better semantic representations present stronger token-token correlations. In the most semantically informative layers of DeepSeek-V3 processing English text, past tokens predict future tokens more accurately than the reverse, highlighting a causal asymmetry in semantic representations.
@@
The presence of semantic layers containing shared information in deep networks is connected to the notion of representation alignment rep_alignment . Two different models can be ‘stitched’ together with Stitching_bansal or without moschella2023 a trainable stitching layer at several different stages of the network. For instance, two vision models can effectively use similar low-level features stitchin0 . Nonetheless, the types of features that we probe for in a semantic layer are instead abstract, encoding the collective relationships between the input constituents, independently of the particular way in which they were expressed in the specific input modality at disposal. The general representations we are probing lie behind the so-called Platonic representation hypothesis, postulating that, as model quality improves, models converge to similar representations platonic_hypothesis . We significantly extend our understanding of the nature of these shared representations with respect to platonic_hypothesis , since for autoregressive language models, the authors take the average across the token axis, reducing the dimensionality of the problem, whereas for vision models, they only consider the “class” token. We address instead where semantic information is encoded within the network, namely across different tokens and different layers. The representation of language has also been studied in Ref. cheng2025emergence using, among other techniques, the Information Imbalance. This work is focused on the analysis of the last-token representation of the same text in different models, and not on multimodal data.
Other earlier works point to deep similarities across architectures and modalities Kornblith:etal:2019 ; Sorscher:etal:2022 ; Maniparambil:etal:2024 . Again, these works do not focus on determining the location and nature of the shared semantic layers. Moreover, they do not include a latest-generation, huge model such as DeepSeek-V3. Our work also relates to studies that probe the extent to which multilingual LLMs produce shared representations across languages (and whether this is due to English acting as a pivot language), e.g.: peng2024conceptspacealignmentmultilingual ; singh-2019-bert ; brinkmann2025largelanguagemodelsshare ; llama_english ; llm_multilingualism . To this research line, we contribute a method to compute similarity that allows us to take long spans of tokens into account and to capture asymmetry between language-specific representations.
@@ 2 The Information Imbalance
Given two feature spaces $A$ and $B$ pertaining to a collection of data points, the II is a measure that quantifies how much information on feature space $B$ is carried by feature space $A$ and vice-versa. For an introduction to the II with simple examples, we refer the reader to Ref. II . In this work, the data points are the high-dimensional representations generated by artificial neural networks. The central idea is to compute the distance between every data point in each feature space separately, obtaining for each point $i$, two sets of distances $d_{i,j}^{A}$ and $d_{i,j}^{B}$, $i,j=1,...,N_{s}$, respectively, where $N_{s}$ is the number of samples. Then, for each data point $i$, all other points $j$ are ranked in each feature space in increasing order, obtaining two sets of integer ranks $r_{i,j}^{A}$ and $r_{i,j}^{B}$. The II between feature spaces $A$ and $B$, denoted by $\Delta(A\rightarrow B)$, is proportional to the average rank of points in space $B$ of the nearest neighbors in space $A$:
$\Delta(A\rightarrow B)=\frac{2}{N_{s}-1}\langle r^{B}|r^{A}=0\rangle\approx\frac{2}{N_{s}-1}\frac{1}{N}\sum_{i,j}\delta_{r_{ij}^{A},0}r_{ij}^{B},$ (1)
where in the last step we estimate the expected value over the dataset, $\langle.\rangle$, as the sample mean, and $\delta_{r_{ij}^{A},0}$ is a Kronecker delta fixing data samples $i,j$ to be first neighbours in space A. Intuitively, if $A$ predicts $B$ to a large degree, the rank in space $B$ of the first neighbors in space $A$ should be small (equal to zero in the case of perfect predictability). Importantly, the II can be asymmetric: if for example $A$ contains more information about $B$ than $B$ about $A$, $\Delta(A\rightarrow B)<\Delta(B\rightarrow A)$. The normalization is chosen such that if feature space A is non-informative about $B$ $\Delta(A\rightarrow B)=1$. While feasible to compute numerically, the II is formally connected to the theory of copula variables II , given that distance ranks are their discrete version, and it provides an upper bound to a conditional mutual information II ; Causality-II . Implementation details are in Supp. Inf., Sec. C.
In this work, we study how the information is spread across the tokens using models where the embedding dimension is of the order of a few thousands and the number of tokens ranges from $40$ to $1024$, bringing to high-dimensional feature spaces of dimension up to $\mathcal{O}(10^{6})$. Thus, to save memory, we binarize the activations through the application of the sign function, a ‘quantization method’ which has been successfully applied in neural network training hubara2018quantized ; guo2018survey ; courbariaux2016binarized ; wang2023bitnet , and in the analysis of representations. For example, it has been shown that it approximately preserves the angles between weights and activations anderson2017high , and recently it was shown to be suitable for calculating the intrinsic dimension of many-token data representations acevedo2025 . It is common practice to clip activations at a fixed quantile and normalize vectors before computing distances, in order to mitigate the effects of outlier dimensions kovaleva-outliers-bert ; removing-outlier-dims and extreme activation values sun2024massive . Binarization implicitly achieves both: it clips all activations to $\pm 1$ and ensures that vectors of the same length have equal norm. For a case in which we can compute the II in full precision, we show in Sec. H of the Supp. Inf. that binarizing and computing distances with the Hamming distance leads to remarkably similar results to those obtained with normalized real-valued activations and $L_{2}$ distances.
@@ 3 Results
We consider sets of features $A_{l}$ and $B_{l}$ in two different languages. The index $l$ labels different layers and ranges from 1 to $L$. If a sentence in the language associated with space $A_{l}$ is tokenized into $T$ tokens, then at each layer $l$, $A_{l}\in\mathds{R}^{T\times E}$, where $E$ denotes the embedding dimension of the language model processing the sentence. What we refer to as the semantic content (or meaning) of a sentence, is a global property, emerging from the interaction of all the symbols present in it. However, in the deep representations the attention mechanism can move information across the tokens, and eventually concentrate it in a specific token, as indeed happens by construction in the last layer of causal models. One of the goals of this work is studying explicitly where the information is stored in the deep layers. This will be addressed in Section 3.1.2. We use human-made translations from Helsinki-NLP/opus_books. Data processing details are in Supp. Inf. D.
@@
Figure 1: Panel a) Information Imbalance (II) from English to Spanish, using representations generated at equal depth from translated sentences of opus_books, as a function of the relative network depth, for DeepSeek-V3 and Llama3.1-8b. A smaller value of the II correspond to higher predictive power. We used the concatenation of the last 2020 tokens for the computation. Panel b) II between equal-layer representations of DeepSeek-V3 processing different translation pairs. For each language pair in the legend, we show the II from the first language to the second. Error bars computed with 5 repetitions of a Jackknife procedure, subsampling $2000$ sentences out of $5000$, are smaller than marker size.
@@
Starting with English-Spanish translations, we compute the information imbalance between $A_{l}$ and $B_{l}$ for every layer in the network, excluding the embedding layer, using the last 20 tokens of each representation (in Sec. 3.1.2 we study the dependence of the II on the number of tokens). Note that $A_{l}$ and $B_{l}$ are generated independently, with the network reading one language at a time, without any specific prompt or instruction. Results are shown in panel a) of Fig. 1. Since the initial layers of LLMs must process the input, the corresponding representations are strongly language-dependent, and hence not highly informative about each other. As we consider deeper layers, we observe that they become progressively more informative about the deep representations in the other language. Since the information shared by both input languages must be semantic in nature, this analysis confirms that the data structure of deep representations seems to encode information in a universal, language-independent manner. For the very last layers of DeepSeek-V3, we observe a sudden rise in Information Imbalance (II), since those layers are generating the output in a specific language, which cannot be universal.
@@
Note that for an ideal dataset of translations, an ideal network separately processing the sentence pairs should generate equivalent representations, namely somewhere in the network the II from one language to the other should be close to zero. Thus, better semantic representations in our experiments have a higher predictability and a lower II. In Fig. 1, for English-Spanish pairs the II of DeepSeek-V3’s representations has a very broad minimum of order 0.2, roughly between relative depths $0.4$ and $0.9$. This number is impressively small, considering that the dimension of the representations which are compared is $20\cdot 7000\sim 150,000$. As a comparison, if the data were generated by Gaussian processes, one would observe an II of 0.2 if approximately 20% of the features were shared between the two representations (see Supp. Inf., Sec. C). Instead, for Llama3.1-8b the minimum II is significantly higher, of order $0.3$, corresponding to approximately 15% shared features. Moreover, the minimum is narrower, around a relative depth of 0.3, suggesting that bigger and better models generate representations of higher semantic content.
@@
We also measured the intrinsic dimensionality of DeepSeek-V3’s representations using the BID estimator acevedo2025 . We find two local maxima of roughly the same height that coincide with the beginning (around 0.4) and the end (around 0.9) of the semantic region observed in Fig. 1, see Supp. Inf., Sec. I. Although these peaks suggest that semantic encoding leverages a higher-dimensional abstract space—echoing findings from brain-similarity studies antonello2024evidence —we defer a deeper investigation of representation dimensionality to future work.
@@
As a sanity check, we show in Supp. Inf. F that performing batch-shuffling on any of the datasets leads to completely uninformative representations (II gives 1 for every layer), since the semantic correspondence between translations is destroyed. Furthermore, in Supp. Inf. G we show that our results are compatible with those obtained using the overlap neighborhood metric from Ref. platonic_hypothesis . In Supp. Inf. E we also show that, if instead of considering many-token representations as we do in Fig. 1, one reduces the dimensionality by averaging all tokens, as done in Ref. platonic_hypothesis , the results are qualitatively compatible, but present significant quantitative differences. For all layers, but most strongly on the first and the last layers, we observe that averaging significantly increases the level of mutual predictability.

### v4

@@ Abstract
Deep neural networks are known to develop similar representations for semantically related data, even when they belong to different domains, such as an image and its description, or the same text in different languages. We present a method for quantitatively investigating this phenomenon by measuring the relative information content of the representations of semantically related data and probing how it is encoded into multiple tokens of large language models (LLMs) and vision transformers. Looking first at how LLMs process pairs of translated sentences, we identify inner “semantic” layers containing the most language-transferable information. We also identify layers encoding semantic information within visual transformers. We show that caption representations in the semantic layers of LLMs predict visual representations of the corresponding images. We observe significant and model-dependent information asymmetries between image and text representations.
@@ 1 Introduction
Current research explores the conjecture that Large Language Models (LLMs) possess a universal language of thought, enabling them to process concepts across different languages [peng2024conceptspacealignmentmultilingual, anthropic_claude, platonic_hypothesis, singh-2019-bert, brinkmann2025largelanguagemodelsshare]. A compelling example comes from Anthropic’s study of the Claude model [anthropic_claude], which shows that the network occasionally reasons within a shared conceptual space across languages. platonic_hypothesis shows that representations of many different LLM/ViTs processing semantically equivalent data are similar to each other, and that this form of alignment is correlated with model size. More generally, representations in deep networks, including but not limited to LLMs, are converging, potentially pointing toward shared underlying conceptual structures.
@@
In this work, we seek to turn the qualitative notion that deep networks develop semantically meaningful shared representations into quantitative measures, identifying where in the network, and to what extent, semantic information is encoded. We can cast the identification of shared semantic content as a problem of relative information: different networks learn distributions defined on different data domains, such as images and captions. Those distributions loosely span a common underlying subspace, associated with shared semantic information. The support of the different distributions does not reduce to this subspace, but spans other subspaces, which are data and architecture specific. To quantify relative information we need a similarity criterion that is both (i) asymmetric, as there may be a partial order relation between models and representations (qualitatively, the subspaces which are data specific might have different dimensions); (ii) computationally efficient for representations of dimension of order $10^{5}-10^{6}$, which is the typical number of neurons in a deep representation. Cross-entropy measures how difficult it is to encode an event from a distribution into another and would be the ideal choice, but estimating it is computationally difficult due to the dimension of the representations. On the other hand, local methods like the average number of nearest neighbors [doimo2020, platonic_hypothesis] fail to capture the asymmetry between spaces. To overcome the limitations, we employ the Information Imbalance [II]: a method that leverages conditional ranking to provide an asymmetric measure of relative mutual predictivity, which has been shown to be an excellent proxy of the cross entropy [Causality-II].
@@
We analyze the representation of translations of the same sentences in different languages, using one of the most powerful open-source language models, DeepSeek-V3 with 671B parameters [deepseekv3], comparing it to the mid-sized Llama3.1-8b LLM [Meta:2024]. Moreover, we analyze the deep representations of pairs of images and of human-generated descriptions, and of pairs of images depicting the same object, with the goal of capturing the alignment between image and caption representations in different architectures and the relative predictive power of images and texts. As visual models, we employ image-gpt-large [pmlr-v119-chen20s] and DinoV2-large [oquabDINOv2LearningRobust2024].
@@
  • We analyze the DeepSeek-V3 language model and measure, layer by layer, the relative information content between representations of sentence pairs that are translations of each other. This reveals a broad region of the network, robust across language pairs, that consistently encodes shared semantic content.
  • We compare different summaries of the data -considering concatenations, mean and last tokens- to quantify their impact on similarity scores. We find that considering several tokens yields significantly stronger similarity, suggesting that semantic information is spread across many tokens.
@@
The presence of semantic layers containing shared information in deep networks is connected to the notion of representation alignment [rep_alignment]. Two different models can be ‘stitched’ together with [Stitching_bansal] or without [moschella2023] a trainable stitching layer at several different stages of the network. For instance, two vision models can effectively use similar low-level features [stitchin0]. Nonetheless, the types of features that we probe for in a semantic layer are instead abstract, encoding the collective relationships between the input constituents, independently of the particular way in which they were expressed in the specific input modality at disposal. The general representations we are probing lie behind the so-called Platonic representation hypothesis, postulating that, as model quality improves, models converge to similar representations [platonic_hypothesis]. We significantly extend our understanding of the nature of these shared representations with respect to [platonic_hypothesis], since for autoregressive language models, the authors take the average across the token axis, reducing the dimensionality of the problem, whereas for vision models, they only consider the “class” token. We address instead where semantic information is encoded within the network, namely across different tokens and different layers. LLM representations have also been studied in cheng2025emergence using, among other techniques, the Information Imbalance. This work is focused on the analysis of the last-token representation of the same text in different models, and not on multimodal data.
Other earlier works point to deep similarities across architectures and modalities [Kornblith:etal:2019, Sorscher:etal:2022, Maniparambil:etal:2024]. Again, these works do not focus on determining the location and nature of the shared semantic layers, nor do they include a latest-generation, large model such as DeepSeek-V3. Our work also relates to studies that probe the extent to which multilingual LLMs produce shared representations across languages (and whether this is due to English acting as a pivot language), e.g. [peng2024conceptspacealignmentmultilingual, singh-2019-bert, brinkmann2025largelanguagemodelsshare, llama_english, llm_multilingualism]. In particular, llama_english analyzes whether multilingual models align non-English inputs by implicitly routing them through English, focusing on logit lens experiments. To this research line, we contribute a method to compute similarity that allows us analyze long spans of tokens and to capture asymmetry between language-specific representations, without the use of the embedding or unembedding matrices. Finally, chughtai2023toy also study universality in learned representations, but while their analysis is purely theoretical and operates in controlled toy settings, our work examines analogous questions using real-world multilingual and multimodal data on actual state-of-the-art LLMs and ViTs.
@@ 2 The Information Imbalance
Figure 1: Left: Information Imbalance $\Delta(X\!\to\!Y)$ and $\Delta(Y\!\to\!X)$, compared with Central Kernel Alignment (CKA), for a synthetic Gaussian construction in which each index $r$ generates a pair $(X_{r},Y_{r})$ via $Y_{r}=B_{r}X_{r}+\varepsilon$, with $X_{r}\sim\mathcal{N}(0,I)$ and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I)$, in $D=10$ dimensions. The matrices $B_{r}\in\mathbb{R}^{D\times D}$ are designed to have monotonically increasing rank—from rank one at r=1r{=}1 to full rank at the final index. Right: Statistical-power benchmark on a high-dimensional Gaussian model. We compute the Information Imbalance and CKA using only a fraction of the $p$ components, ranging from small subsets to the full vector for $p=10^{3}$ and $p=10^{5}$. In both figures, we report the standard error computed by averaging over ten jackknife repetitions.
Large language models encode information in high-dimensional spaces, transforming representations across layers to perform a task. The Platonic representation hypothesis [platonic_hypothesis] suggests that –for relevant model sizes– representations converge to similar neighboring structures regardless of (i) the task of the model and (ii) the specific encoding of the information. The emerging regularity of these structures underpins the idea that meaning drives their organization, with models loosely acting as continuous mappings from a hidden manifold in which semantically similar ideas are close to each other. We can thus cast the convergence to a universal representation as a mutual information problem. If we model pairs of representations as random variables with a joint probability distribution, for instance an image and its caption, we can measure the information lost by replacing the joint distribution $p(x,y)$ with the product of the marginals $p(x)p(y)$. If the distributions are independent, observing captions is not informative of the density of images. If representations converge, knowing a caption is highly informative and the joint distribution is not approximated well by the marginals.
Mutual information alone, however, is not sufficient to compare different models. The Platonic hypothesis argues that convergence towards a universal set of representations depends on model size. We therefore need to determine (i) whether the effect correlates with the specific encoding of the information (images, captions, …); (ii) whether architectural differences matter; and (iii) how convergence changes with model size. From an information-theoretic standpoint, we need an asymmetric measure that quantifies how much uncertainty over the representation in $Y$ (e.g., captions) remains once we observe the representation in $X$ (e.g., images). In practice, estimating information-theoretical measures which could quantify such asymmetry is difficult in the high-dimensional spaces produced by modern models. For this reason, we turn to the Information Imbalance [II, Causality-II], a rank-based measure that tests whether neighbourhoods in one representation space predict neighbourhoods in another: if representations converge, the nearest neighbours of a point in $Y$ should also be nearest neighbours of the corresponding point in $X$ in the other architecture.
Formally, the Information Imbalance compares the neighborhood structures of two representation spaces $X$ and $Y$. Given representations $\{\mathbf{z}_{i}^{X}\}_{i=1}^{N}$ and $\{\mathbf{z}_{i}^{Y}\}_{i=1}^{N}$, we compute all pairwise distances and rank points $j\neq i$ in each space by increasing distance, obtaining ranks $r_{i,j}^{X}$ and $r_{i,j}^{Y}$. The Information Imbalance from $X$ to $Y$ is the normalized average rank in $Y$ of the nearest neighbor of $i$ in $X$:
$\Delta(X\rightarrow Y)=\frac{2}{N-1}\,\frac{1}{N}\sum_{i=1}^{N}r_{i,\mathrm{NN}(i)}^{Y},$
where $\mathrm{NN}(i)$ denotes the index of the closest point to $i$ in space $X$, i.e. $r_{i,\mathrm{NN}(i)}^{X}=1$. If neighborhoods in $X$ predict neighborhoods in $Y$, these ranks are small and $\Delta(X\rightarrow Y)$ is close to zero; if $X$ carries no information about $Y$, the expected rank is uniform and $\Delta(X\rightarrow Y)\approx 1$. The asymmetry $\Delta(X\rightarrow Y)\neq\Delta(Y\rightarrow X)$ quantifies directional predictability across models or modalities. In Fig. 1 (left), we show that the information imbalance captures both the strength and the direction of the relative information across layers in a controlled synthetic setting, highlighting how asymmetry emerges when information is compressed or degraded. We compute the Information Imbalance $\Delta(X\!\to\!Y)$ and $\Delta(Y\!\to\!X)$ for a synthetic Gaussian construction in which each index $r$ generates a pair $(X_{r},Y_{r})$ via $Y_{r}=B_{r}X_{r}+\varepsilon$, with $X_{r}\sim\mathcal{N}(0,I)$ and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I)$, in $D=10$ dimensions. The $B_{r}$s are defined by first choosing a target rank $r\in\{1,\dots,D\}$ and then sampling independent Gaussian factors $U_{r}\sim\mathcal{N}(0,1)^{D\times r}$ and $V_{r}\sim\mathcal{N}(0,1)^{r\times D}$, with the linear map given by $B_{r}=U_{r}V_{r}$, which enforces a transformation of rank $r$. High–rank settings preserve most of the structure in $X_{r}$, whereas low–rank settings collapse information into a low-dimensional, noisy subspace. The information-imbalance curves reveal the resulting directional improved predictivity, showing that $\Delta(X\!\to\!Y)$ is lower than $\Delta(Y\!\to\!X)$. For comparison, we also report CKA, which tracks the overall loss of alignment between $X_{r}$ and $Y_{r}$ but, being symmetric, cannot expose the asymmetry captured by Information Imbalance. A complementary statistical-power benchmark is shown in Fig. 1 (right), where we compute the Information Imbalance between a gaussian vectors with $p$ components and a vector containing only given a fraction of them, for cases $p=10^{3}$ and $p=10^{5}$. For a $10^{5}$-dimensional Gaussian, CKA saturates to 1 even when only a small fraction of the components is observed, while the Information Imbalance retains discriminative resolution up to a much larger fraction of shared features. This illustrative example echoes the findings of platonic_hypothesis, that showed their neighborhood overlap was preferable to CKA in high dimensions.
Our representations live in extremely high-dimensional spaces, with embedding dimensions of a few thousand across 404010241024 tokens, leading to feature spaces of size up to $\mathcal{O}(10^{6})$. A central difficulty is that it is not clear a priori which summary of these token-level representations should be used: many studies rely on the last token or on the average, but these choices may discard information. One advantage of the Information Imbalance is that it scales to high dimensions, allowing us to concatenate all tokens and systematically compare different summaries without committing to a specific one. To reduce memory usage and stabilise distance computations, we binarize activations using the sign function—a quantization method widely used in neural network training [hubara2018quantized, guo2018survey, courbariaux2016binarized, wang2023bitnet] and in representation analysis, where it approximately preserves angles [anderson2017high] and has proved suitable for studying the intrinsic dimension of many-token representations [acevedo2025]. Standard practice in representation analysis involves clipping activations and normalising vectors to mitigate outlier dimensions [kovaleva-outliers-bert, removing-outlier-dims, sun2024massive]; binarization implicitly performs both steps by clipping values to $\pm 1$ and enforcing equal norms. In Sec. G of the Supp. Inf., we show that binarized and floating-point activations yield essentially identical information-imbalance curves, indicating that our results depend on the geometric arrangement of points—relative distances and neighbourhoods—rather than on the precise numerical values of the activations.
@@ 3 Results
We consider sets of features $A_{l}$ and $B_{l}$ in two different languages. The index $l$ labels different layers and ranges from 1 to $L$. If a sentence in the language associated with space $A_{l}$ is tokenized into $T$ tokens, then at each layer $l$, $A_{l}\in\mathds{R}^{T\times E}$, where $E$ denotes the embedding dimension of the language model processing the sentence. What we refer to as the semantic content (or meaning) of a sentence, is a global property, emerging from the interaction of all the symbols present in it. However, in the deep representations the attention mechanism can move information across the tokens, and eventually concentrate it in a specific token, as indeed happens by construction in the last layer of causal models. One of the goals of this work is studying explicitly where the information is stored in the deep layers. This will be addressed in Section 3.1.2. We use human-made translations from Helsinki-NLP/opus_books. Data processing details are in Supp. Inf. C.
@@
Figure 2: Panel a) Information Imbalance from English to Spanish, using representations generated at equal depth from translated sentences of opus books, as a function of the relative network depth, for DeepSeek-V3, BERT-multilingual and Llama3 with 1, 3, and 8 billion parameters. A smaller value of the Information Imbalance correspond to higher predictive power. We used the concatenation of the last 2020 tokens for the computation. Panel b) Information Imbalance between equal-layer representations of DeepSeek-V3 processing different translation pairs. For each language pair in the legend, we show the Information Imbalance from the first language to the second. The Jackknife error bars (five repetitions subsampling 2,000 of 5,000 sentences) are too small to show; the marker sizes act as an upper bound on these errors.
Starting with English-Spanish translations, we compute the Information Imbalance between $A_{l}$ and $B_{l}$ for every layer in the network, excluding the embedding layer, using the last 20 tokens of each representation (in Sec. 3.1.2 we study the dependence of the Information Imbalance on the number of tokens). Note that $A_{l}$ and $B_{l}$ are generated independently, with the network reading one language at a time, without any specific prompt or instruction. We also verified that our conclusions are robust to standard preprocessing and similarity metrics, such as clipping-and-normalization and the neighborhood overlap of platonic_hypothesis, as detailed in Supp. Inf. F.
@@
Results are shown in panel a) of Fig. 2. Since the initial layers of LLMs must process the input, we expect the corresponding representations to be strongly language-dependent, and hence not highly informative about each other. As we consider deeper layers, we observe that they become progressively more informative about the deep representations in the other language. Since the information shared by both input languages must be semantic in nature, this analysis confirms that the data structure of deep representations seems to encode information in a universal, language-independent manner. In the very last layers of DeepSeek-V3, we see an apparent jump in Information Imbalance. Part of this effect is due to the fact that we report only a subsample of layers for computational reasons, which makes the transition appear more abrupt. Beyond this sampling artefact, the increase is expected: the final layers are driven by next-token prediction and therefore encode strongly language-specific representations, which cannot be universal.
@@
Note that -under the Platonic representation hypothesis- for an ideal dataset of translations, an ideal network separately processing the sentence pairs should generate equivalent representations, namely somewhere in the network the Information Imbalance from one language to the other should be close to zero. Thus, we define semantic quality in terms of mutual predictability, which should manifest as a lower Information Imbalance. In Fig. 2, for English-Spanish pairs the Information Imbalance of DeepSeek-V3’s representations has a very broad minimum of order 0.2, roughly between relative depths $0.4$ and $0.9$. This number is impressively small, considering that the dimension of the representations which are compared is $20\cdot 7000\sim 150,000$. As a comparison, if the data were generated by Gaussian processes, one would observe an Information Imbalance of 0.2 if approximately 20% of the features were shared between the two representations (see Fig. 1). Instead, for Llama3.1-8b the minimum Information Imbalance is significantly higher, of order $0.3$, corresponding to approximately 15% shared features. Moreover, the minimum is narrower, around a relative depth of 0.3, suggesting that bigger and better models generate representations of higher semantic content.
Finally, to account for different model sizes, we performed the same experiments on two additional Llama models, observing that the minimum Information Imbalance moves downward and gets slightly displaced toward earlier layers. Notably, this effect only takes place in inner layers, whereas the II measured in both initial and final layers does not seem to scale down with model size. To account for a difference in training objective, we also added a multilingual BERT model with 0.2B parameters, also showing a qualitatively similar behavior.
@@
We also measure the intrinsic dimensionality of DeepSeek-V3’s representations using the BID estimator [acevedo2025]. We find two local maxima of roughly the same height that coincide with the beginning (around 0.4) and the end (around 0.9) of the semantic region observed in Fig. 2, see Supp. Inf., Sec. H. Although these peaks suggest that semantic encoding leverages a higher-dimensional abstract space—echoing findings from brain-similarity studies [antonello2024evidence]—we defer a deeper investigation of representation dimensionality to future work.
@@
As a null hypothesis, we show in Supp. Inf. E that performing batch-shuffling on any of the datasets leads to completely uninformative representations (Information Imbalance gives 1 for every layer), since the semantic correspondence between translations is destroyed. Furthermore, in Supp. Inf. F we show that our results are compatible with those obtained using the overlap neighborhood metric from platonic_hypothesis. In Supp. Inf. D we also show that, if instead of considering many-token representations as we do in Fig. 2, one reduces the dimensionality by averaging all tokens, as done in platonic_hypothesis, the results are qualitatively compatible, but present significant quantitative differences. For all layers, but most strongly on the first and the last layers, we observe that averaging significantly increases the level of mutual predictability.